### v2

@@ Abstract
Zichen Zhu, Hao Tang, Yansi Li, Dingye Liu, Hongshen Xu, Kunyao Lan, Danyang Zhang, Yixuan Jiang, Hao Zhou, Chenrun Wang, Situo Zhang, Liangtai Sun, Yixiao Wang, Yuheng Sun, Lu Chen [^1] Corresponding authors., Kai Yu [^1] Corresponding authors. X-LANCE Lab, Department of Computer Science and Engineering MoE Key Lab of Artificial Intelligence, SJTU AI Institute Shanghai Jiao Tong University, Shanghai, China {JamesZhutheThird,chenlusz,kai.yu}@sjtu.edu.cn
@@ 1 Introduction
Multimodal large language models (MLLMs) have seen significant advancements in recent years, supported by vast multimodal datasets. These models Hu et al. (2024); Liu et al. (2024a); Ye et al. (2024, 2023); Chen et al. (2024b); Sun et al. (2024a); Liu et al. (2023); Dai et al. (2023); Chen et al. (2023); Zhu et al. (2024); Yao et al. (2024); OpenAI (2023); Team (2024) excel in tasks such as Chain-of-Thought (CoT) reasoning Wei et al. (2022), In-Context Learning (ICL) Brown et al. (2020), and various applications Wang et al. (2024c); Wang and Zhao (2023); Chen et al. (2024a); Liu et al. (2024b); Pan et al. (2024); Ge et al. (2024); Wu et al. (2024); Lee et al. (2024b); Qian et al. (2024b, a). Their capabilities have also enabled new MLLM-based agents for real-world tasks Li et al. (2017, 2019); Sun et al. (2022); Zhu et al. (2023); Zhan and Zhang (2023); Zhang et al. (2023a, 2024a); Nong et al. (2024); Ma et al. (2024); Wang et al. (2024b, a).
{CJK}
UTF8gbsn
@@
However, MLLMs face significant challenges when addressing complex GUI interactions and facing diverse user demands in real-world scenarios, particularly on devices such as smartphones Zhang et al. (2024b) and computers Cao et al. (2024); Xie et al. (2024). On the one hand, GUI environments are highly diverse and pose different action spaces across different apps and pages. For instance, the number and position of clickable icons can vary greatly across pages; some pages require text input, while others involve scrollable elements. Such variability makes proactive task planning hardly adapt to the real environment contexts and thus become infeasible to complete. On the other hand, the action executor can also lack capabilities enough to achieve it, even given a feasible task plan. In all these cases, agents with trivial or static planning Zheng et al. (2024); Zhang et al. (2024a); Nong et al. (2024); Ma et al. (2024); Xing et al. (2024a) will fail to align with the environment contexts and action executor’s capacity and thus can fail the whole task easily caused by failure of a single sub-task. Furthermore, existing MLLM-based GUI agents Zhang et al. (2023b, a); Wang et al. (2024b, a) often lack a powerful and comprehensive memory to face the need for dynamic planning at various levels and diverse user demands. These problems hinder the design of a practical mobile assistant.
@@
To address these challenges, we propose MobA, a novel MLLM-based mobile assistant system with an adaptive planning module that dynamically adjusts task plans according to the execution results. As proactive planning often fails to accurately determine the actions required in a specific application or page or to align with the action executor’s capacity, MobA leverages reflection mechanism to recover task execution from failed sub-plans by reassessing goals or breaking tasks into more fine-grained sub-goals. To better support adaptive planning with various sub-goal granularity, a multifaceted memory module providing hierarchical memory support is proposed. We also introduce MobBench, a diverse dataset for complex mobile interactions, and demonstrate MobA’s effectiveness on MobBench and AndroidArena Xing et al. (2024b), showing its capability to handle dynamic GUI environments.
@@
  • We introduce MobBench, a diverse dataset for complex mobile interactions, and validate the effectiveness of our approach through extensive experiments on MobBench and AndroidArena.
Figure 1: The Illustration of adaptive planning and multifaceted memory structure. There are 4 cases in adaptive planning: (a) Plan reflection failure, indicating the goal needs to be decomposed. (b) In execution reflection failure, the goal needs to be decomposed. (c) Execution reflection failure, the goal needs to be refined. (d) Goal complete.
@@ 2 The MobA System
{CJK}
UTF8gbsn {CJK}UTF8gbsn {CJK}UTF8gbsn MobA comprises a Global Agent (GA) and a Local Agent (LA). The Global Agent consists of a Plan Module and a Reflection Module. The Plan Module interprets the user’s command (➀) and resolves the task into several easier and clearer sub-tasks adaptively with the help of experiences in the multifaceted Memory Module (➁), while the Reflection Module will review if the decomposition is feasible and sub-goals are achievable. Then, under the direction of a specific sub-goal (➂), the Local Agent will leverage the experiences in the Memory Module (➃), predict the concrete actions, and directly control the device (➄). After LA’s execution, the Reflection Module will reflect if the current sub-task has been completed (➅) and the Plan Module can revise the plan accordingly. The Memory Module can also be updated after the invocation of the Plan Module and Local Agent (➆) to improve MobA’s performance through execution. To ameliorate the performance at the early stage of the memory, it can also be initialized with warm-up of some basic expert experiences. Finally, MobA can generate a response to the user regarding the result of task execution (➇). The remaining parts of this section will elaborate on the proposed adaptive plan module and multifaceted memory module.
@@ 3 Experiments
To comprehensively compare MobA with other GUI agents in handling complex user instructions and executing GUI interactions on mobile devices, we evaluate them using a real-life scenario test set called MobBench. Additionally, we assess our method using the widely adopted mobile benchmark, Android Arena.
@@ 3.1 The MobBench Test Set
Compared with several similar task sets mentioned in other papers Zhang et al. (2023a); Wang et al. (2024b, a); Zhang et al. (2024a); Lee et al. (2024a), which only get a score when it finishes the task, we assign several milestone scores for sub-tasks in MobBench. This allows for a more precise process assessment, in the cases where the task is partially finished. We also include a detailed preparation instruction for tasks when a more justice and stable start is needed.
@@ 3.2 Metrics
Task Type # Tasks Avg. MS Avg. Steps EE
Easy 10 1.0 4.3 4.30
Medium 10 2.2 7.3 3.32
Hard 10 4.1 15.2 3.71
Indirect 10 2.8 9.4 3.36
Cross-App 10 3.1 10.8 3.48
Overall 50 2.7 9.4 3.56
Table 1: Average scores and expert execution steps for different task types of MobBench.
@@ 3.3 Setups
Mobile Agent (v2) Wang et al. (2024b, a) uses only visual information from screenshots as inputs. Target elements are selected with the guidance of OCR and CLIP Radford et al. (2021) models.
@@ 3.4 Results and Analysis
Human 50/50 133 3.56
GPT-4o + Human 49/50 130 (97.7%) 3.82 (107.2%)
AppAgent 6/50 35 (28.6%) 4.43 (124.4%)
MobileAgent (v2) 17/50 63 (48.9%) 4.84 (136.0%)
MobA w/o M & P 13/50 52 (39.1%) 4.42 (124.2%)
MobA w/o P 15/50 65 (48.9%) 4.17 (117.1%)
MobA w/o M 22/50 72 (54.1%) 3.81 (106.9%)
MobA 28/50 88 (66.2%) 3.44 (96.7%)
@@ 3.5 Results on Android Arena
We also performed evaluations on Android Arena Xing et al. (2024b), comprising 157 single-app tasks and 21 cross-app tasks. As shown in Table 3, MobA achieves success rates (SR) of 0.783 on single-app tasks and 0.714 on cross-app tasks, outperforming GPT-4 by 2.4% and 14.3%, respectively. The notable improvement in cross-app tasks is attributed to MobA’s subtask decomposition capability, which enables better app-switching decisions during tasks requiring more steps. Additionally, MobA’s reflection module encourages exploration, reducing repetitive actions and improving task success rates.
@@ 5 Conclusion and Future Works
This paper presented MobA, an innovative Mobile phone Assistant system empowered by MLLMs. Utilizing a two-level agent structure, comprising a Global Agent and a Local Agent, MobA effectively understands user commands, plans tasks, and executes actions. The combination of Memory and Plan Modules enhances its ability to learn from previous interactions, improving efficiency and accuracy. Our evaluations demonstrated that MobA surpasses existing mobile assistants in handling complex tasks, leveraging multi-level memory, task decomposition, and action-validation mechanisms. These features enable precise task execution even with intricate or indirect commands. Future work will focus on improving the performance on image-only scenarios where the view hierarchy is unattainable, deploying an end-side model on mobile phones for faster response and secured privacy. We hope MobA illustrates the potential of MLLMs-empowered mobile assistants and provides valuable insights for future works.
@@ Appendix A Several Useful Links
Appendix B Related Work
B.1 LLM Agents
The development of intelligent agents has been significantly influenced by the advancements in large language models (LLMs) and multimodal large language models (MLLMs). LLM-based agents leverage the autonomy, reactivity, proactiveness, and social ability of these models to perceive external environments and make decisions Xi et al. (2023). Emerging abilities, such as chain-of-thought (CoT) reasoning Wei et al. (2022); Wang et al. (2023b); Zhang et al. (2023d) and in-context learning (ICL) Brown et al. (2020); Min et al. (2022). Recent studies have explored LLM-based approaches for reflection Yao et al. (2023); Madaan et al. (2023); Shinn et al. (2023); Xu et al. (2024b), planning Sun et al. (2024b); Qian et al. (2024c); Huang et al. (2024), and memory mechanisms Zhang et al. (2024d, c); Li et al. (2023); Maharana et al. (2024); Lan et al. (2024); Zhang et al. (2023c).
B.2 GUI Agents
B.2.1 Traditional GUI Agents
Controlling graphical user interface (GUI) screens based on user commands is a complex task that involves both GUI understanding and command interpretation. Early approaches to GUI agents focused on embedding and modular systems. For example, several agents Li et al. (2017, 2019) combined natural language and programming demonstrations, allowing users to define tasks via descriptions and demonstrations. This method relied on text and image matching for script-based control of the interface. Traditional GUI agents were largely limited by their reliance on pre-defined rules and manual programming. These agents were effective within controlled environments but struggled with dynamic, real-world GUI contexts due to their lack of flexibility and adaptability. They required specific scripts or rules for each task, making them less robust when handling the diverse and evolving nature of real-world applications.
B.2.2 Advancements with Multimodal Pretrain Models
The advent of multimodal pretraining models Bai et al. (2021); Li et al. (2021b); Li and Li (2023); He et al. (2021); Li et al. (2021a); Wang et al. (2021); Fu et al. (2024) for GUI understanding marked a significant shift in the development of GUI agents. Pretrained agents Sun et al. (2022); Zhu et al. (2023); Zhan and Zhang (2023); Xu et al. (2024a) integrated multimodal information, such as dialogue history, screenshots, and action history, through pretraining. Unlike earlier methods that relied on rigid scripts, these end-to-end models adopted a more human-like approach to interacting with interfaces, enhancing their efficiency in information retrieval and task execution by mapping visual observations and text commands directly into actions.
B.2.3 MLLM-Empowered GUI Agents
The integration of MLLMs in GUI agents has introduced new opportunities to further enhance their capabilities. With the rise of larger scale models, GUI agents Zhang et al. (2023a, 2024a); Lee et al. (2024a) began to leverage advanced reasoning and decision-making processes. These models utilized structural information provided in the view hierarchy (VH) to annotate and locate UI elements, guiding a sequence of atomic actions to achieve specific goals. VH-only agents Wen et al. (2024) depend on the structural information to reason and make decisions, which greatly lowers the cost of inference making it suitable for deployment on the device. Image-only agents Wang et al. (2024b, a); Gao et al. (2024); Yan et al. (2023), which employs optical character recognition (OCR), CLIP Radford et al. (2021) module, and object detection methods to identify operation targets. This image-only approach is particularly effective when the view hierarchy is inaccessible or noisy, but it may also encounter challenges, e.g. opening a target application by clicking when names are hidden, or logos vary across screens.
Appendix C View hierarchy processing
@@
Appendix D Action Space
@@ 3. Suitable message,
Appendix F Detailed Results Comparison
@@
F.1 Human is more adaptive and robust to screen interactions

### v3

@@ Abstract
Zichen Zhu, Hao Tang, Yansi Li, Dingye Liu, Hongshen Xu, Kunyao Lan, Danyang Zhang, Yixuan Jiang, Hao Zhou, Chenrun Wang, Situo Zhang, Liangtai Sun, Yixiao Wang, Yuheng Sun, Lu Chen [^1] Corresponding authors are Lu Chen and Kai Yu., Kai Yu [^1] Corresponding authors are Lu Chen and Kai Yu. X-LANCE Lab, Department of Computer Science and Engineering MoE Key Lab of Artificial Intelligence, SJTU AI Institute Shanghai Jiao Tong University, Shanghai, China {JamesZhutheThird,chenlusz,kai.yu}@sjtu.edu.cn
@@ 1 Introduction
Multimodal large language models (MLLMs) have seen significant advancements in recent years, supported by vast multimodal datasets. These models Hu et al. (2024); Liu et al. (2024a); Ye et al. (2024, 2023); Chen et al. (2024b); Sun et al. (2024a); Liu et al. (2023); Dai et al. (2023); Chen et al. (2023); Zhu et al. (2024); Yao et al. (2024); OpenAI (2023); Team (2024) excel in tasks such as Chain-of-Thought (CoT) reasoning Wei et al. (2022), In-Context Learning (ICL) Brown et al. (2020), and various applications Wang et al. (2024b); Wang and Zhao (2023); Chen et al. (2024a); Liu et al. (2024b); Pan et al. (2024); Ge et al. (2024); Wu et al. (2024); Lee et al. (2024b); Qian et al. (2024b, a). Their capabilities have also enabled new MLLM-based agents for real-world tasks Li et al. (2017, 2019); Sun et al. (2022); Zhu et al. (2023); Zhang and Zhang (2024); Zhang et al. (2023a, 2024a); Nong et al. (2024); Ma et al. (2024); Wang et al. (2024a, 2025).
@@
However, MLLMs face significant challenges when addressing complex GUI interactions and facing diverse user demands in real-world scenarios, particularly on devices such as smartphones Zhang et al. (2024b) and computers Cao et al. (2024); Xie et al. (2024). On the one hand, GUI environments are highly diverse and pose different action spaces across different apps and pages. For instance, the number and position of clickable icons can vary greatly across pages; some pages require text input, while others involve scrollable elements. Such variability makes proactive task planning hardly adapt to the real environment contexts and thus become infeasible to complete. On the other hand, the action executor can also lack capabilities enough to achieve it, even given a feasible task plan. In all these cases, agents with trivial or static planning Zheng et al. (2024); Zhang et al. (2024a); Nong et al. (2024); Ma et al. (2024); Xing et al. (2024) will fail to align with the environment contexts and action executor’s capacity and thus can fail the whole task easily caused by failure of a single sub-task. Furthermore, existing MLLM-based GUI agents Zhang et al. (2023b, a); Wang et al. (2024a, 2025) often lack a powerful and comprehensive memory to face the need for dynamic planning at various levels and diverse user demands. These problems hinder the design of a practical mobile assistant.
Figure 1: The Illustration of adaptive planning and multifaceted memory structure. There are 4 cases in adaptive planning: (a) Plan reflection failure, the goal needs to be decomposed. (b) In execution reflection failure, the goal needs to be decomposed. (c) Execution reflection failure, the goal needs to be refined. (d) Goal complete.
@@
To address these challenges, we propose MobA, a novel MLLM-based mobile assistant system with an adaptive planning module that dynamically adjusts task plans according to the execution results. As proactive planning often fails to accurately determine the actions required in a specific application or page or to align with the action executor’s capacity, MobA leverages reflection mechanism to recover task execution from failed sub-plans by reassessing goals or breaking tasks into more fine-grained sub-goals. To better support adaptive planning with various sub-goal granularity, a multifaceted memory module providing hierarchical memory support is proposed. We also introduce MobBench, a diverse dataset for complex mobile interactions, and demonstrate MobA’s effectiveness on MobBench and AndroidArena Xing et al. (2024), showing its capability to handle dynamic GUI environments.
@@
  • We introduce MobBench, a diverse dataset for complex mobile interactions, and validate the effectiveness of our approach through extensive experiments on two datasets.
@@ 2 The MobA System
The system overview of MobA is shown in Figure 2. MobA comprises a Global Agent (GA) and a Local Agent (LA). The Global Agent consists of a Plan Module and a Reflection Module. The Plan Module interprets the user’s command (➀) and resolves the task into several easier and clearer sub-tasks adaptively with the help of experiences in the multifaceted Memory Module (➁), while the Reflection Module will review if the decomposition is feasible and sub-goals are achievable. Then, under the direction of a specific sub-goal (➂), the Local Agent will leverage the experiences in the Memory Module (➃), predict the concrete actions, and directly control the device (➄). After LA’s execution, the Reflection Module will reflect if the current sub-task has been completed (➅) and the Plan Module can revise the plan accordingly. The Memory Module can also be updated after the invocation of the Plan Module and Local Agent (➆) to improve MobA’s performance through execution. To ameliorate the performance at the early stage of the memory, it can also be initialized with a warm-up of some basic expert experiences. Finally, MobA can generate a response to the user regarding the result of task execution (➇). The remaining parts of this section will elaborate on the proposed adaptive plan module and multifaceted memory module.
@@ 3 Experiments
To comprehensively compare MobA with other GUI agents in handling complex user instructions and executing GUI interactions on mobile devices, we evaluate them using a real-life scenario test set called MobBench. Additionally, we assess our method using a widely adopted mobile benchmark, Android Arena.
@@ 3.1 The MobBench Test Set
Compared with several similar task sets mentioned in other papers Zhang et al. (2023a); Wang et al. (2024a, 2025); Zhang et al. (2024a); Lee et al. (2024a), which only get a score when it finishes the task, we assign several milestone scores for sub-tasks in MobBench. This allows for a more precise process assessment, in the cases where the task is partially finished. We also include a detailed preparation instruction for tasks when a more reproducible, fair, and stable start is needed.
@@ 3.2 Metrics
Task Type # Tasks # MS Avg. Steps EE
Easy 10 10 4.3 4.30
Medium 10 23 7.3 3.17
Hard 10 41 15.2 3.71
Indirect 10 28 9.4 3.36
Cross-App 10 31 10.8 3.48
Overall 50 133 9.4 3.53
Table 1: Milestone scores and expert execution steps for different task types of MobBench.
@@ 3.3 Setups
Mobile Agent (v2) Wang et al. (2024a, 2025) uses only visual information from screens as inputs. Target elements are selected with the guidance of OCR and CLIP Radford et al. (2021) modules.
@@ 3.4 Results and Analysis
Human 50/50 133 3.53
GPT-4o + Human 49/50 131 (98.5%) 3.82 (108.2%)
AppAgent 6/50 35 (26.3%) 4.43 (125.5%)
MobileAgent (v2) 17/50 63 (47.4%) 4.84 (137.1%)
MobA w/o M & P 13/50 52 (39.1%) 4.42 (125.2%)
MobA w/o P 15/50 65 (48.9%) 4.17 (118.1%)
MobA w/o M 22/50 72 (54.1%) 3.81 (107.9%)
MobA 28/50 88 (66.2%) 3.44 (97.5%)
@@ 3.5 Results on Android Arena
We also performed evaluations on Android Arena Xing et al. (2024), comprising 157 single-app tasks and 21 cross-app tasks. As shown in Table 3, MobA achieves success rates (SR) of 0.783 on single-app tasks and 0.714 on cross-app tasks, outperforming GPT-4 by 2.4% and 14.3%, respectively. The notable improvement in cross-app tasks is attributed to MobA’s subtask decomposition capability, which enables better app-switching decisions during tasks requiring more steps. Additionally, MobA’s reflection module encourages exploration, reducing repetitive actions and improving task success rates.
@@ 4 Case Study
5 Related Work
5.1 LLM Agents
The advancements in M/LLMs have significantly influenced the development of agents. LLM-based agents leverage the autonomy, reactivity, proactiveness, and social ability of these models to perceive external environments and make decisions Xi et al. (2023). Emerging abilities, such as CoT reasoning Wei et al. (2022); Wang et al. (2023b); Zhang et al. (2023c) and in-context learning Brown et al. (2020); Min et al. (2022). Recent studies have explored LLM-based approaches for reflection Yao et al. (2023); Madaan et al. (2023); Shinn et al. (2023); Xu et al. , planning Sun et al. (2024b); Qian et al. (2024c); Huang et al. (2024), and memory mechanisms Zhang et al. (2024d, c); Li et al. (2023); Maharana et al. (2024); Lan et al. (2024).
5.2 GUI Agents
5.2.1 Traditional GUI Agents
Controlling GUI screens based on user commands is a complex task that involves both GUI understanding and command interpretation. Early approaches to GUI agents focused on embedding and modular systems. For example, several agents Li et al. (2017, 2019) combined natural language and programming demonstrations, allowing users to define tasks via descriptions and demonstrations. This method relied on text and image matching for script-based control of the interface. Traditional GUI agents were largely limited by their reliance on pre-defined rules and manual programming. These agents were effective within controlled environments but struggled with dynamic, real-world GUI contexts due to their lack of flexibility and adaptability. They required specific scripts or rules for each task, making them less robust when handling the diverse and evolving nature of real-world applications.
5.2.2 Advancements with Multimodal Pretrain Models
The advent of multimodal pretraining models Bai et al. (2021); Li et al. (2021b); Li and Li (2023); He et al. (2021); Li et al. (2021a); Wang et al. (2021); Fu et al. (2024) for GUI understanding marked a significant shift in the development of GUI agents. Pretrained agents Sun et al. (2022); Zhu et al. (2023); Zhang and Zhang (2024); Xu et al. (2024) integrated multimodal information, such as dialogue history, screenshots, and action history, through pretraining. Unlike earlier methods that relied on rigid scripts, these end-to-end models adopted a more human-like approach to interacting with interfaces, enhancing their efficiency in information retrieval and task execution by mapping visual observations and text commands directly into actions.
5.2.3 MLLM-Empowered GUI Agents
The integration of MLLMs in GUI agents has introduced new opportunities to further enhance their capabilities. With the rise of larger scale models, GUI agents Zhang et al. (2023a, 2024a); Lee et al. (2024a) began to leverage advanced reasoning and decision-making processes. These models utilized structural information provided in the view hierarchy (VH) to annotate and locate UI elements, guiding a sequence of atomic actions to achieve specific goals. VH-only agents Wen et al. (2024) depend on the structural information to reason and make decisions, which greatly lowers the cost of inference making it suitable for deployment on the device. Image-only agents Wang et al. (2024a, 2025); Gao et al. (2024); Yan et al. (2023), which employs OCR, CLIP Radford et al. (2021) module, and object detection methods to identify operation targets. This image-only approach is particularly effective when the view hierarchy is inaccessible or noisy, but it may also encounter challenges, e.g. opening a target application by clicking when names are hidden, or logos vary across screens.
@@ 6 Conclusion and Future Works
This paper presented MobA, an innovative Mobile phone Assistant system empowered by MLLMs. Utilizing a two-level agent structure, comprising a Global Agent and a Local Agent, MobA effectively understands user commands, plans tasks, and executes actions. The combination of Memory and Plan Modules enhances its ability to learn from previous interactions, improving efficiency and accuracy. Our evaluations demonstrated that MobA surpasses existing mobile assistants in handling complex tasks, leveraging multi-level memory, task decomposition, and action-validation mechanisms. These features enable precise task execution even with intricate or indirect commands. Future work will focus on improving the performance on image-only scenarios where the view hierarchy is unattainable, deploying an end-side model on mobile phones for faster response and secured privacy. We will continue to expand MobBench by adding more popular applications from different regions and languages. We hope MobA illustrates the potential of MLLMs-empowered mobile assistants and provides valuable insights for future works.
Acknowledgment
This work is funded by the China NSFC Projects (92370206, 62120106006, U23B2057) and Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102).
@@ Appendix A Several Useful Links
Appendix B View hierarchy processing
@@
Appendix C Action Space
@@ 3. Suitable message,
Appendix E Detailed Results Comparison
@@
E.1 Human is more adaptive and robust to screen interactions